OpenStudy (anonymous):

Trig let sin t=a , cos t=b , and tan t=c. write the expression in terms of a, b, and c. 5 cos (-t) - cos t

OpenStudy (perl):

Hint : cos(-t) = cos (t)

OpenStudy (anonymous):

why do you say that?

Nnesha (nnesha):

bec it's a rule \[\huge\rm sin( - \theta) = - \sin \theta \] ^^^odd \[\huge\rm cos (-\theta)= \cos \theta\] ^^even function

OpenStudy (anonymous):

okay so… 5 cos (t) - cos t ?

OpenStudy (perl):

replace cost t with b

OpenStudy (anonymous):

5b-b ?

OpenStudy (perl):

simplify that

OpenStudy (anonymous):

4b ?